40m QIL Cryo_Lab CTN SUS_Lab CAML OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m elog  Not logged in ELOG logo
Message ID: 6981     Entry time: Tue Jul 17 18:00:58 2012
Author: Masha 
Type: Update 
Category: PEM 
Subject: STS 

Den and investigated the STS-1 problem (which is currently plugged into ADC channels 13, 14, and 15, which correspond to the STS-3 channels in dtt). It turns out that I had plugged in the power to the monitor in the host box rather than the remote. The X, Y, and Z readout is currently approaching a mean of zero, and I will let it continue to do so overnight (pressing auto-zero as necessary). Attached is a plot of the coherence with GUR 1, and the time-domain signals.

Attachment 1: Screenshot.png  66 kB  | Hide | Hide all
Screenshot.png
ELOG V3.1.3-